An uncle is twice as old as his nephew, and 8 years ago he was three times as old as his nephew. Find their present ages.
Answer by lwsshak3(11628)

An uncle is twice as old as his nephew, and 8 years ago he was three times as old as his nephew. Find their present ages.
***
At present:
let x=nephew's age
2x=uncle's age
..
8 years ago:
x-8=nephew's age
2x-8=uncle's age
2x-8=3(x-8) (8 years ago he was three times as old as his nephew.)
2x-8=3x-24
x=16
nephew's age=16
uncle's age=32